OpenStudy (anonymous):

Factor completely: 10x2 + 2x − 8

OpenStudy (jkristia):

can you see how to start ?. Is there a common factor in all 3 terms ?

OpenStudy (anonymous):

the choices are: 2(5x − 1)(x + 4) 2(5x − 4)(x + 1) 2(5x + 2)(x − 2) 2(5x − 2)(x + 2)

OpenStudy (anonymous):

it is 2.

OpenStudy (anonymous):

then it looks like 5x^2 + x - 4

OpenStudy (anonymous):

is it the third one? 2(5x + 2)(x − 2)?

OpenStudy (jkristia):

sorry - I got distracted...

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

yes or no please... :/

OpenStudy (jkristia):

when I have to factor a quadratic like that I do it mostly by trial and error (I don't remember the cross multiply thing) You know that one term must start with 5x, so (5x + a?)(x - b?) since you 5x*b (+-) a?*x = 1, then obviously b cannot 4, so you have (5x +(-)4)(x +(-)b?) now you know 4 * b = 4, so b - 1. Now it is a matter finding the correct signed for (+)(-). Sorry - that was probably not very helpful, but those are the steps I usually do when factoring. If I get really stuck I use the quadratic formula as a last resort.

OpenStudy (jkristia):

try distribute (5x + 2)(x − 2) = 5x^2 - 10x + 2x - 4 = 5x^2 - 8x - 4 so that is not the answer

OpenStudy (anonymous):

ok then you just switch the signs. is it the last one?

OpenStudy (jkristia):

try distribute (5x − 2)(x + 2) 5x^2 + 10x - 2x - 4 5x^2 + 8x - 4 so no :)

OpenStudy (anonymous):

but then you have to multiply the 5x by 2

OpenStudy (jkristia):

2(5x^2 + x - 4) so ignore the 2 for now and look for something where (5x +- A)(x +- B) = 5x^2 + x - 4 so 5x(B) - A(x) = 1, (1x) and (A)(B) = -4

OpenStudy (jkristia):

I'm not really good at explaining this - sorry about that.

OpenStudy (anonymous):

can you tell me what one it is? I am totally confused...

OpenStudy (jkristia):

2(5x-4)(x+1) = 2(5x^2 + 5x - 4x - 4) = 2(5x^2 + 1x - 4) = 10x^2 + 2x - 8

OpenStudy (jkristia):

you see that?

OpenStudy (anonymous):

ok